I just can't seem to get this. Can you help me solve this equation? Thanks
x/8 +x/12 = -4

Answer by checkley77(12844)

x/8 +x/12 = -4
(3x+2x)/24=-4
5x/24=-4
5x=-4*24
5x=-96
x=-96/5
x=-19.2 answer.
Proof:
-19.2/8-19.2/12=-4
-2.4-1.6=-4
-4=-4